# Simulations of gravitational collapse in null coordinates IV: evolving through the event horizon, with an application to the spherical charged scalar field

### v1

@@ VI.1 Previous work
By contrast, in GellesPretorius25 , the spherical Maxwell-charged scalar system is evolved on a fixed RN spacetime, in order to investigate event horizon instabilities. That work uses double-null coordinates, compactified to both past and future null infinity (but finite across the past and future event horizon) as in MurReaTan13 . It uses the electromagnetic gauge $\nabla_{a}(R^{-2}A^{a})=0$, or $A_{u,v}+A_{v,u}=0$ in double-null coordinates. (This is in contrast to our electromagnetic gauge $A_{x}=0$, which in double-null coordinates becomes $A_{v}=0$.)
@@ VI.2 Physical setup
Our algorithm also allows us to fix the right spacetime gauge as $R_{,x}=R_{,x*}$ (so that $x$ is a linear funtion of $R$), but for brevity we do not present results here. When there is no scalar field on the right boundary, Bondi and affine right gauge are equivalent anyway. Morever, in our discretisation they are equivalent without discretisation error.
We can continue the spacetime gauge into the numerical domain with any choice of $B(u,x)$ (as long as $B=0$ on the left boundary), but again for brevity we only present double-null gauge $B(u,x)=0$, so that $x$ is an ingoing null coordinate everywhere. We use the timestep $\Delta u_{0}=C_{0}\Delta x$, with $C=0.1$.
We shall restrict to scenarios where $C_{*}<1$, so that the corner 2-sphere is untrapped. Then $R_{,x*}>0$ gives us $R_{,u*}<0$, and the ingoing Raychaudhuri equation (49) gives us $R_{,u}<0$ on the entire left boundary. We also restrict to the case where the right boundary is untrapped, so that $R_{,x}>0$ there.
@@ VI.3 Prompt collapse
Figure 8: Spacetime diagram of the prompt collapse scenario, showing the regular centre and numerical domain in bold, the singularity in bold-dashed, and the event horizon as a thin line.
Figure 9: Spacetime diagram of collapse from data on a null cone with regular centre, with the lines as in Fig. 8. For suitable data, this could be the same spacetime as in Fig. 8.
Consider a null rectangle with $R_{*}>0$, $R>0$ on the left boundary, $M_{*}=Q_{*}=0$, zero scalar field on the left boundary, and no trapped surface on the right boundary. Then the triangular region bounded by the regular centre $R=0$, $u=u_{0}$ and $v=v_{0}$ is a piece of Minkowski spacetime, and $u=u_{0}$ can be extended to the left to $R=0$ and to the right to $R=\infty$ as a lightcone with a regular vertex and no trapped surface.
If the data on $u=0$ are sufficiently strong that a MOTS first forms at $(v_{1},u_{\text{FMOTS}})$ for some $u_{0}<u_{\text{FMOTS}}<u_{1}$, then there must be an event horizon at some $u=u_{\text{EH}}$ with $u_{0}<u_{\text{EH}}\leq u_{\text{FMOTS}}$. We have in effect simulated gravitational collapse from initial data on a null cone with regular centre, while actually solving the Einstein equations only on a null rectangle, and thus avoiding the complications of a regular centre. The null rectangle may then be truncated by a curvature singularity “at the top”. Figs. 8 and 9 show gravitational collapse in the null rectangle and regular centre settings.
@@
As an example of prompt collapse, we set ${\cal M}_{*}=Q_{*}=0$, $R_{*}=2$, and fix the corner gauge as $G_{*}=R_{,x*}=1$. With no scalar field on the left boundary, $R_{,u}=-1/2$ there, so the left boundary intersects $R=0$ at $u=4$. On the right boundary we set $\psi$ to be a Gaussian with centre at $x_{c}=0.5$, width $d=0.1$, frequency $\omega=50$ and amplitude $9.5\cdot 10^{-3}$. We set $q=50$. From now on, without loss of generality, we let $u_{0}=0$ and $v_{0}=v$, and we identify $v=0$ with $x=0$. The event horizon forms just after $u=3.3$. In the eG and fe formulations we evolve to $u=3.9$, just before singularity excision starts, and in the eR formulation at $u=3.3$, just before the event horizon forms.
@@ VI.4 Schwarzschild and extremal Reissner-Nordström
We continue to extremal RN. We set ${\cal M}_{*}=Q_{*}=1$, so that the horizon radius is again $r_{+}=1$, and we again set $R_{*}=2$. $M$ is now not constant, and in particular at the corner $M_{*}=3/4$, so that $C_{*}=3/8$. We again fix the gauge as $G_{*}=R_{,x*}=1$. This gives us $R_{,u}=-1/8$ on the left boundary, so that the event horizon is crossed at $u=8$, and the left boundary runs into the singularity at $u=16$.
@@
We evolve with $N_{x}=100\cdot 2^{n}$, with n=06n=0...6, up to $u=7.9$ in the eR formulation, and up to $u=14$ in the eG and fe formulations. We see perfect self-convergence and convergence between formulations in the highest resolutions, but convergence in the lower resolutions begins to break down between $u=13$ and $u=14$ due to the appearance of sharp gradients in all variables near the left boundary.
@@ VI.5 Perturbed extremal Reissner-Nordström
Not unexpectedly, this particular attempt fails. Indeed for all $u>1.7$, already well before the putative horizon is crossed, $Q\leq{\cal M}\leq 1$ everywhere on the domain $0\leq x\leq 1$ (with equality only at the left boundary, where the solution is still electrovacuum extremal RN). The event horizon is crossed at $u\simeq 9$ and has radius $R\simeq 0.9$ (and hence Hawking mass $M\simeq 0.45$), augmented mass ${\cal M}\simeq 0.996$ and charge $Q\simeq 0.991$, so the final black hole is slightly subcritical.
However, while $R$ on our numerical domain $0\leq x\leq 1$ levels off in $x$ at $u\simeq 9$, it returns to being an increasing function of $x$ afterwards, at least up to $u=14$, when we stop the simulation. Although this behaviour is characteristic of an extremal black hole (which has no trapped surfaces in the interior), we know in this case that the black hole is clearly subcritical. Hence the fact that we do not see $R_{,x}<0$ or the spacelike singularity $R=0$ on our numerical domain must be due to its limited size.
We evolve with $N_{x}=100\cdot 2^{n}$, now with n=06n=0...6. We need the higher resolutions to see clean self-convergence and convergence between resolutions. The simulation is more demanding than that of unperturbed eRN because $\phi$ becomes more rapidly oscillating in $x$: in the initial data, $\phi$ has a about two complex oscillations within a Gaussian envelope, while at $u=14$ there are about 30 complex oscillations in roughly the same envelope.
@@
Once again we see good convergence within formulations and between formulations at the highest resolution. At $u=14$, the complex scalar field is rapidly oscillating in $x$, with about 20 phase oscillations between $x=0$ and $x=0.1$, with an exponentially decaying envelope, and this is only just resolved at the two highest resolutions.
@@ VII Conclusions
This series of papers investigates the usefulness of “Bondi-like” null coordinates, in which surfaces of constant retarded time $u$ are null cones and the coordinate lines of constant $(u,\theta,\varphi)$ are their generators. The remaining gauge freedom is how points on each generator are labelled by the radial coordinate $x$.
We have identified three generic formulations: the metric coefficient $R$ is evolved and the metric coefficient $G$ is found by solving the (outgoing) Raychaudhuri equation, or the other way around, or both $R$ and $G$ are evolved freely (and the Raychaudhuri equation is imposed as a constraint only at $u=0$). In these three, any labelling of points on a null generator by the radial coordinate $x$ is possible. Dynamically, this radial gauge choice is controlled by freely choosing $B$. There are also two particular formulations each tied to a specific gauge: the Bondi formulation where $R=x$, and the affine formulation where $G=1$.
@@
To present the five different formulations pedagogically, and with a focus on regularity at the central worldline $R=0$, in this paper we have restricted attention to spherical symmetry. We will explore the eG formulation in axisymmetry in a future paper, but we notice already that the generalisation beyond spherical symmetry and/or to higher dimensions is conceptually straightforward.
As a nontrivial test-bed in spherical symmetry, we have implemented the spherical Einstein-Maxwell-charged scalar system, allowing for initial data either on a single outgoing null cone $u=0$ emerging from a regular centre $R=0$, or on two intersecting null cones $u=u_{0}$ and $v=v_{0}$.
In both setups we have shown second-order self-convergence in all three formulations and second-order convergence between all three formulations. With a regular centre, the convergence is pointwise except near the centre, where it still holds in any norm, such as the root-mean-square or maximum norm. The fact that even the fe formulation converges without any tweaks at the centre came as a pleasant surprise. With the left boundary an ingoing null cone, the convergence is perfectly pointwise everywhere.
As a physics testbed of collapse from regular data on an outgoing null cone, we have investigated critical collapse of a charged scalar field. This had previously been investigated theoretically in GundlachMartin96 , and numerically in HodPiran97 ; Petryk05 , but here we show the first accurate plots of the universal fine structure in the mass and charge scaling laws. Surprisingly, we found that being able to evaluate the mass and charge on a marginally trapped surface makes a qualititative difference to our results, compared to evaluating them on an almost-trapped surface, even one with Hawking compactness as large as $C=0.999$. We also found initial evidence that the fine-structure of the charge scaling law may not be universal. We will return to this elsewhere.
@@
However, the main motivation for introducing the two formulations in which $G$ is evolved is that in them we can evolve through event and apparent horizons. In Paper II, both our formulation (we evolved $R$) and our choice of gauge (“local shifted Bondi gauge”) did not allow us to do this in axisymmetric scalar collapse. This in turn meant that we could not bisect arbitrarily close to the collapse threshold in highly-nonspherical one-parameter families of initial data. We plan to apply our eG formulation with a suitable radiable gauge to this problem.
@@
As a physics testbed of the setup with initial data on two intersecting null cones, we have simulated gravitational collapse with a Minkowski interior, the exact Schwarzschild and extreme Reissner-Nordström solutions (which are non-trivial in our gauge), and extreme Reissner-Nordström perturbed by a complex scalar field on either $u=u_{0}$ or $v=v_{0}$.
@@
We plan a study of the solution space of the spherical Einstein-Maxwell-charged scalar system, and in particular an investigation of the “extremal critical collapse conjecture for the spherical Einstein-charged scalar field system KehleUnger24 , see also KehleUnger22 . We anticipate that both types of numerical domain will be required for this investigation.
@@ Appendix A The Reissner-Nordström metric in null coordinates
and the real constants ${\cal M}_{0}>0$ and $Q_{0}$ are the usual mass and charge parameters of the RN metric, normalised such that $|Q_{0}|={\cal M}_{0}$ represents an extremal black hole. The event and Cauchy horizon are at
@@
By definition the Hawking mass obeys $r=2M$ on the event horizon. $M=r_{+}/2$ at the event horizon is also the irreducible mass of the black hole. In the special case of an extremal black hole, $r_{+}=r_{-}=2M={\cal M}=Q$ on the event horizon.
@@
we have $u(\bar{u}_{0})=u_{0}$, $v(\bar{v}_{0})=v_{0}$, $R_{,u}=-1/a$ along $v=v_{0}$ and $R_{,v}=1/b$ along $u=u_{0}$. Inverting, we obtain
@@
and hence $G(u_{0},v)=G(u,v_{0})=G_{*}$. This is, in implicit form, the exterior metric of RN in the coordinates we use in our null-rectangle tests, with $(R_{,v})_{*}=1/b$, $R_{,u*}=-1/a$, and
$G_{*}={1\over ab\,\bar{G}(\bar{u}_{0},\bar{v}_{0})}.$ (108)
@@ Appendix B Regular centre
The Minkowski metric in Bondi coordinates, with $U$ normalised to be proper time at the centre, is
@@
From elementary flatness, all these conditions at the centre must also hold in an arbitary curved spherical spacetime with a regular centre $R=x=0$ and $u$ proper time at the centre. Beyond spherical symmetry, in spherical coordinates, they will also hold at the regular “central” worldline $R=x=0$ as long as it is geodesic, see Appendix III.E.2 of Paper I for this subtlety.

### v2

@@ VI.1 Previous work
By contrast, in GellesPretorius25 , the spherical Maxwell-charged scalar system is evolved on a fixed RN spacetime, in order to investigate event horizon instabilities. That work uses double-null coordinates, compactified to both past and future null infinity (but finite across the past and future event horizon) as in MurReaTan13 . It uses the electromagnetic gauge $\nabla_{a}(R^{-2}A^{a})=0$, or $A_{u,v}+A_{v,u}=0$ in double-null coordinates, in contrast to our electromagnetic gauge $A_{x}=0$, which in double-null coordinates becomes $A_{v}=0$.
@@ VI.2 Physical setup
Our algorithm also allows us to fix the right spacetime gauge as $R_{,x}=R_{,x*}$, but for brevity we do not present results here. When there is no scalar field on the right boundary, Bondi and affine gauge are equivalent there anyway. Morever, in our discretisation they are equivalent without discretisation error.
We can continue the spacetime gauge into the numerical domain with any choice of $B(u,x)$ (as long as $B=0$ on the left boundary), but again for brevity we only present double-null gauge $B(u,x)=0$, so that $x$ is an ingoing null coordinate everywhere, and the numerical domain is the full domain of dependence of the data (a null rectangle). We use the timestep $\Delta u_{0}=C_{0}\Delta x$, with $C_{0}=0.1$.
We shall restrict to scenarios where $C_{*}<1$, so that the corner 2-sphere is untrapped. Then $R_{,x*}>0$ gives us $R_{,u*}<0$, and the ingoing Raychaudhuri equation (49) gives us $R_{,u}<0$ on the entire left boundary. We also restrict to the case where the entire right boundary is untrapped, so that $R_{,x}>0$ there.
@@ VI.3 Prompt collapse
Figure 9: Spacetime diagram illustrating collapse from initial data on an outgoing null cone, with singularity excision. The regular centre and numerical domain are shown as bold lines, the singularity as bold dashed, lines of constant $u$ as thin solid, and lines of constant $x$ as thin dashed. The lines of constant $x$ illustrate sdn gauge, with $x=0$ the timelike regular centre $R=0$, $x=x_{0}$ ingoing null, and $x=x_{\text{max}}$ future spacelike. Initial data can be imposed on the outgoing null cone (with regular centre) $u=0$, symbolised here by a bell curve.
Figure 10: Spacetime diagram illustrating the null rectangle setup, with initial data imposed on the ingoing null cone $x=0$ and/or the outgoing null cone $u=0$. The lines of constant $u$ (thin solid) are outgoing null again, as in Fig. 9, but the lines of constant $x$ (thin dashed) are all ingoing null, illustrating double-null gauge.
Figure 11: Spacetime diagram illustrating what we have called “prompt collapse”. Free initial data are imposed on $u=0$, but the data on $x=0$ are now specifically for Minkowski spacetime, so that the solution on the numerical domain can be continued into the triangle bounded by $R=0$, $x=0$ and $u=0$ (extended as thin solid lines) as a part of Minkowski spacetime.
Consider a null rectangle with $R_{*}>0$, $R>0$ on the left boundary, $M_{*}=Q_{*}=0$, zero scalar field on the left boundary, and no trapped surface on the right boundary. Then the triangular region bounded by the regular centre $R=0$, $u=u_{0}$ and $v=v_{0}$ is a piece of Minkowski spacetime, and $u=u_{0}$ can be extended to the left to $R=0$, with vanishing scalar field.
If the data on $u=0$ are sufficiently strong that a MOTS first forms at $(v_{1},u_{\text{FMOTS}})$ for some $u_{0}<u_{\text{FMOTS}}<u_{1}$, then there must be an event horizon at some $u=u_{\text{EH}}$ with $u_{0}<u_{\text{EH}}\leq u_{\text{FMOTS}}$. We have in effect simulated gravitational collapse from initial data on a null cone with regular centre, while actually solving the Einstein equations only on a null rectangle, and thus avoiding the complications of a regular centre. The null rectangle may then be truncated by a curvature singularity “at the top”. Figs. 9 and 10 show gravitational collapse from data on a single null cone, and from data on two intersecting null cones. Fig. 11 illustrates the prompt collapse scenario. In each case, we illustrate excision of a future spacelike singularity from the numerical domain.
@@
As an example of prompt collapse, we set ${\cal M}_{*}=Q_{*}=0$, $R_{*}=2$, and fix the corner gauge as $G_{*}=R_{,x*}=1$. With no scalar field on the left boundary, $R_{,u}=R_{,u*}=-1/2$ there, so the left boundary intersects $R=0$ at $u=4$. On the right boundary we set $\psi$ to be a Gaussian with centre at $x_{c}=0.5$, width $d=0.1$, frequency $\omega=50$ and amplitude $9.5\cdot 10^{-3}$. We set $q=50$ and $x_{\text{max}}=1$. From now on, without loss of generality, we let $u_{0}=0$ and $v_{0}=v$, and we identify $v=0$ with $x=0$. The event horizon forms just after $u=3.3$. In the eG and fe formulations we evolve to $u=3.9$, just before singularity excision starts, and in the eR formulation at $u=3.3$, just before the event horizon forms.
@@ VI.4 Schwarzschild and extremal Reissner-Nordström
We continue to extremal RN. We set ${\cal M}_{*}=Q_{*}=1$, so that the horizon radius is again $r_{+}=1$, and we again set $R_{*}=2$. $M$ is now not constant, and in particular at the corner $M_{*}=3/4$, so that $C_{*}=3/4$. We again fix the gauge as $G_{*}=R_{,x*}=1$. This gives us $R_{,u}=R_{,u*}=-1/8$ on the left boundary, so that the event horizon is crossed at $u=8$, and the left boundary runs into the singularity at $u=16$. We set $x_{\text{max}}=1$.
@@
We evolve with $N_{x}=100\cdot 2^{n}$, with n=06n=0...6, up to $u=7.9$ in the eR formulation, and up to $u=14$ in the eG and fe formulations. We see perfect self-convergence and convergence between formulations in the highest resolutions, but convergence in the lower resolutions begins to break down between $u=13$ and $u=14$, due to the appearance of sharp gradients in all variables near the left boundary.
@@ VI.5 Perturbed extremal Reissner-Nordström
Not unexpectedly, this particular attempt fails. Indeed for all $u>1.7$, already well before the putative horizon is crossed, $Q\leq{\cal M}\leq 1$ everywhere on the domain $0\leq x\leq 1$ (with equality only at the left boundary). The event horizon is crossed at $u\simeq 7.4$, that is, earlier than in the unperturbed eRN solution and has radius $R\simeq 1.08$ (and hence Hawking mass $M\simeq 0.54$), augmented mass ${\cal M}\simeq 0.997$ and charge $Q\simeq 0.993$, so the final black hole is slightly subcritical.
$R$ on our numerical domain $0\leq x\leq 1$ levels off in $x$ at $u\simeq 7.4$, indicating the event horizon, then has a maximum in $x$ between $u\simeq 7.4$ and $u\simeq 8.8$, and then returns to be an increasing function of $x$ at least up to $u=14$, when we stop the simulation. This is illustrated in Fig. 12. That we do not see $R_{,x}<0$ for sufficiently large $x$, for all $u>u_{\text{horizon}}$, is probably due to the limited size of our numerical domain. We will investigate this elsewhere.
Figure 12: Plot of the area radius $R(u,x,)$ in the eRN solution perturbed by a complex Gaussian on the right boundary $u=0$. We plot against $x$, at $u=0$, $7.2$, $7.4$, $7.6$ and $14$ (from top to bottom). The lower plot shows only $u=7.2$, $7.4$ and $7.6$ for clarity, demonstrating that the event horizon is crossed at $u\simeq 7.4$.
We evolve with $N_{x}=100\cdot 2^{n}$, with n=06n=0...6. We need the higher resolutions to see clean self-convergence and convergence between resolutions. The simulation is more demanding than that of unperturbed eRN because $\phi$ becomes rapidly oscillating in $x$. This is illustrated in Fig. 13.
Figure 13: Plot of the real part $\psi(u,x,)$ of the complex scalar field in the eRN solution perturbed by a complex Gaussian on the right boundary $u=0$. We plot against $x$, at $u=0$ (the simpler graph) and at $u=14$ (the highly oscillating graph). The imaginary part $\chi$ looks similar, with a phase shifting indicating that the rapid oscillation is in the complex phase. The numerical domain is $0\leq x\leq 1$, but we plot only the range where $\psi$ is large.
@@
Once again we see good convergence within formulations and between formulations at the highest resolution. At $u=14$, the complex scalar field is rapidly oscillating in $x$, with about 16 phase oscillations between $x=0$ and $x=0.1$, with an exponentially decaying envelope, and this is only just resolved at the two highest resolutions.
@@ VII Conclusions
This series of papers investigates the usefulness of single null coordinates, in which surfaces of constant retarded time $u$ are null cones and the coordinate lines of constant $(u,\theta,\varphi)$ are their generators. The remaining gauge freedom is how points on each generator are labelled by the radial coordinate $x$.
We have identified three generic formulations: the metric coefficient $R$ is evolved and the metric coefficient $G$ is found by solving the (outgoing) Raychaudhuri equation, or the other way around, or both $R$ and $G$ are evolved freely (and the Raychaudhuri equation is imposed as a constraint only at $u=0$). In these three, any labelling of points on a null generator by the radial coordinate $x$ is possible. Dynamically, this radial gauge choice is controlled by freely choosing $B$. There are also two particular formulations each tied to a specific gauge: the Bondi formulation where $R=x$, and the affine formulation where $G=1$. These are maximally constrained, in the sense that only physical degrees of freedom are evolved.
@@
The main motivation for introducing the eG and fe formulations, in which $G$ is evolved, is that in them we can evolve through event and apparent horizons. In Paper II, both our choice of the eR formulation and our choice of lsB gauge did not allow us to do this in axisymmetric scalar collapse. This in turn meant that we could not bisect arbitrarily close to the collapse threshold in highly-nonspherical one-parameter families of initial data. In the eG and fe formulations, we can now evolve through horizons, while still using $B$ to make the coordinate adjust to the approximate self-similarity observed in type-II critical collapse. We plan to apply our eG formulation with a suitable radiable gauge to this problem.
To present the new formulations pedagogically, and with a focus on regularity at the central worldline $R=0$, in this paper we have restricted attention to spherical symmetry. The generalisation beyond spherical symmetry and/or to higher dimensions is conceptually straightforward, and will be given elsewhere. As an example in spherical symmetry that is also of interest in its own right, we have implemented the spherical Einstein-Maxwell-charged scalar system. We allow for initial data either on a single outgoing null cone $u=0$ emerging from a regular centre $R=0$, or on two intersecting null cones $u=u_{0}$ and $v=v_{0}$.
In both setups we have shown second-order self-convergence in all three formulations and second-order convergence between all three formulations. With a regular centre, the convergence is pointwise except near the centre, where it still holds in any function norm. The fact that even the fe formulation converges without any tweaks at the centre came as a pleasant surprise. With the left boundary an ingoing null cone, the convergence is perfectly pointwise everywhere.
As a physics testbed of collapse with data on an outgoing null cone, we have investigated spherical type-II critical collapse of a charged scalar field. This had previously been investigated theoretically in GundlachMartin96 , and numerically in HodPiran97 ; Petryk05 , but here we have presented the first accurate plots of the universal fine structure in the mass and charge scaling laws. We have also presented some initial evidence for non-universality of the charge scaling fine-structure.
@@
As testbeds for the setup with initial data on two intersecting null cones, we have simulated gravitational collapse with a Minkowski interior, the exact Schwarzschild and extreme Reissner-Nordström solutions (which are non-trivial in our gauge), and extreme Reissner-Nordström perturbed by a complex scalar field on either $u=u_{0}$ or $v=v_{0}$.
@@
We plan a study of the solution space of the spherical Einstein-Maxwell-charged scalar system, and in particular an investigation of the “extremal critical collapse conjecture” for the spherical Einstein-charged scalar system KehleUnger24 , see also KehleUnger22 . We anticipate that both types of numerical domain will be required for this investigation.
@@ Appendix A The Reissner-Nordström metric in null coordinates
[…]
@@
By definition the Hawking mass obeys $r=2M$ on the event horizon. $M=r_{+}/2$ at the event horizon is also the irreducible mass of the black hole. In the special case of an extremal black hole, $r_{+}=r_{-}=2M={\cal M}=Q$.
@@
we have $u(\bar{u}_{0})=u_{0}$, $v(\bar{v}_{0})=v_{0}$, $R_{,u}(u,v_{0})=R_{,u*}:=-1/a$ and $R_{,v}(u_{0},v)=(R_{,v})_{*}:=1/b$. Inverting (99,100), we obtain
@@
and hence $G(u_{0},v)=G(u,v_{0})=G_{*}$, where
$G_{*}:={1\over ab\,\bar{G}(\bar{u}_{0},\bar{v}_{0})}.$ (108)
This is, in implicit form, the exterior metric of RN in the coordinates we use in our null-rectangle tests.
@@ Appendix B Regular centre
The Minkowski metric in Bondi coordinates, with $U$ the retarded time normalised to be proper time at the centre, is
@@
From elementary flatness, these conditions at the centre must also hold in an arbitary curved spherical spacetime with a regular centre $R=x=0$ and $u$ proper time at the centre. Beyond spherical symmetry, in spherical coordinates, they will also hold at the regular worldline $R=x=0$ as long as it is geodesic, see Appendix III.E.2 of Paper I.